OpenStudy (anonymous):

Solve each equation below on the interval 0 ≤ x < 2π. a. sin 2x - √2sin x = 0 b. sin x + 2tan x = 0 c. 2sin² x + 3cos x – 3 = 0

OpenStudy (anonymous):

a)factorıse sın x \[2 \sin x \cos x=\sqrt 2 \sin x\] b) \[\sin x+2\frac{\sin x}{\cos x}=0\] c) \[\text{use } \sin ^2x=1-\cos^2x\] then factorıse